# PDDFormer: Pairwise Distance Distribution Graph Transformer for Crystal Material Property Prediction

### v4

@@ Abstract
The crystal structure can be simplified as a periodic point set repeating across the entire three-dimensional space along an underlying lattice. Traditionally, methods for representing crystals rely on descriptors like lattice parameters, symmetry, and space groups to characterize the structure. However, in reality, atoms in material always vibrate above absolute zero, causing continuous fluctuations in their positions. This dynamic behavior disrupts the underlying periodicity of the lattice, making crystal graphs based on static lattice parameters and conventional descriptors discontinuous under even slight perturbations. To this end, chemists proposed the Pairwise Distance Distribution (PDD) method, which has been used to distinguish all periodic structures in the world’s largest real materials collection, the Cambridge Structural Database. However, achieving the completeness of PDD requires defining a large number of neighboring atoms, resulting in high computational costs. Moreover, it does not account for atomic information, making it challenging to directly apply PDD to crystal material property prediction tasks. To address these challenges, we propose the atom-Weighted Pairwise Distance Distribution (WPDD) and Unit cell Pairwise Distance Distribution (UPDD) for the first time, incorporating them into the construction of multi-edge crystal graphs. Based on this, we further developed WPDDFormer and UPDDFormer, graph transformer architecture constructed using WPDD and UPDD crystal graphs. We demonstrate that this method maintains the continuity and completeness of crystal graphs even under slight perturbations in atomic positions. Moreover, by modeling PDD as global information and integrating it into matrix-based message passing, we significantly reduced computational costs. Comprehensive evaluation results show that WPDDFormer achieves state-of-the-art predictive accuracy across tasks on benchmark datasets such as the Materials Project and JARVIS-DFT.
@@ 1 Introduction
Crystals are solids with a regular geometric shape formed by atoms, ions, or molecules arranged periodically in space during the crystallization process. Their structure is typically described using repeating unit cells and lattice vectors. However, this method of description brings a fundamental challenge: the same crystal structure can be represented by different unit cells and lattice vectors, as shown in Figure 1. Additionally, in real-world scenarios, the experimental coordinates of unit cells and atoms are inevitably affected by atomic vibrations and measurement noise. These subtle disturbances can lead to discontinuous changes in any simplified unit cell Kurlin (2024), resulting in numerous different unit cells for a given crystal structure, as shown in Figure 1, thereby introducing ambiguity in the representation of crystal data Widdowson and Kurlin (2022). Currently, many graph neural networks Batzner et al. (2022); Yan et al. (2022, 2024a, 2024b) typically use unit cell parameters, simplified cell parameters, symmetry, and space groups to represent the periodic structure of crystals. However, these features are either non-invariant or discontinuous Zwart et al. (2008) invariants, leaving the issue of ambiguity in crystal data unresolved Patterson (1944); Widdowson et al. (2022); Groom et al. (2016); Bartók et al. (2013); Wassermann et al. (2010); Ahmad et al. (2018).
The continuous and complete invariant—Pairwise Distance Distribution (PDD)—proposed by Widdowson and Kurlin (2022) addresses the ambiguity in crystal data representation by distinguishing all periodic structures in the world’s largest real material collection, the Cambridge Structural Database. To achieve completeness, PDD requires a predetermined number of sufficient neighbors, which is computationally expensive and difficult to directly apply for predicting crystal properties Balasingham et al. (2022). Balasingham et al. (2024) employed distance distribution graphs (DDGs) based on PDD to predict the properties of crystal materials, but they did not achieve satisfactory performance (only slightly better than CGCNN), and although this approach reduced computational costs, it compromised the completeness of PDD. In contrast, crystal graph representations based on multi-edge crystal graphs and unit cell parameters Taniai et al. (2024); Yan et al. (2024a) achieve completeness, more accurately characterizing crystal structures and achieving state-of-the-art performance in crystal material property prediction tasks. However, the use of unit cell parameters leads to discontinuities in the crystal graphs.
Since PDD does not account for atomic types, it is challenging to use it directly for effective crystal property prediction. To better represent crystal structures, we first introduce atom-Weighted PDD (WPDD) and intra-Unit cell PDD (UPDD). Furthermore, we integrate WPDD and UPDD into the construction of multi-edge crystal graphs and propose the PDD Graph Transformer (including WPDDFormer and UPDDFormer) based on the transformer architecture. We model WPDD as global information and incorporate it into matrix-based message passing, significantly reducing computational costs (as shown in Table 3). Finally, we employ the Earth Mover’s Distance (EMD) Rubner et al. (2000) to assess the continuity of crystal graphs, demonstrating that WPDD crystal graphs constructed using only Euclidean distances maintain continuity and completeness under slight atomic position perturbations, providing a more accurate depiction of real crystal structures. Ablation experiments show the crucial role of (W/U)PDD in constructing crystal graphs. Through comprehensive evaluations, our method achieves state-of-the-art predictive accuracy across various tasks in the Materials Project Chen et al. (2019) and JARVIS-DFT Choudhary et al. (2020) datasets. This advancement highlights the effectiveness of WPDDFormer in bridging the gap between traditional crystal descriptors and dynamic atomic behavior, leading to more accurate and reliable predictions in materials science.
@@
Figure 1: Illustrations of different unit cell and lattice representations of the same crystal structure. The blue area in the figure represents possible unit cell structures. Figure (a) shows several possible choices among the infinitely many unit cells for the same crystal structure in the undisturbed case. Figure (b) illustrates that for almost any perturbation, the symmetry group and any reduced unit cell (with minimal volume) will undergo discontinuous changes. This figure is derived from the original PDD Figures 1 and 2.
@@ 2 Preliminaries
In this section, we introduce the definitions of crystal structures, PDD, isometric crystal graphs, and the continuity and geometric completeness of crystal graphs. Additionally, we provide in Appendix B the definition and proof of the unique geometric constraints of crystals.
@@ 2.1 The structure of crystals
By selecting an appropriate structural unit, the entire crystal structure can be viewed as the periodic repetition of this unit in space. This property, where atoms within a crystal repeat in three-dimensional space according to a specific pattern, is called periodicity, with the smallest repeatable structural unit being the unit cell. The unit cell can be defined as $\mathcal{U}=\left(\mathcal{X},\mathcal{P}\right)$, where $\mathcal{X}$ and $\mathcal{Z}$ can be represented in matrix form. Typically, $\mathcal{X}=\left[x_{1},x_{2}\cdots x_{n-1},x_{n}\right]^{T}\in\mathbb{R}^{n\times 1}$, where $n$ represents the number of atoms and ${x_{i}}\in\mathbb{R}^{1}$ represents the atomic type of atom ${i}$ in the unit cell. $\mathcal{P}=\left[p_{1},p_{2}\cdots p_{n-1},p_{n}\right]^{T}\in\mathbb{R}^{n\times 3}$ is the atomic position matrix, where ${p_{i}}\in\mathbb{R}^{3}$ represents the Cartesian coordinates of the atom ${i}$ in the unit cell in 3D space. The lattice vectors $\mathcal{L}=\left[l_{1},l_{2},l_{3}\right]^{T}\in\mathbb{R}^{3\times 3}$ can reflect the way the unit cell repeats in three directions to map the periodic crystal structure. Therefore, in 3D space, the infinite crystal structure $\mathcal{S}$ can be represented as $\left(\mathcal{U},\mathcal{L}\right)$.
2.2 Continuity and geometric completeness of crystal graphs
@@
Definition 1: Pointwise Distance Distribution. For the infinite crystal structure $\mathcal{S}=\left(\mathcal{U},\mathcal{L}\right)$ mentioned in Section 2.1, fix a neighbor count $k\geq 1$. For each point $x_{i}$ in the unit cell $\mathcal{U}$, let $d_{i1}\leq\dots\leq d_{ik}$ be the Euclidean distances from $\textbf{p}_{i}$ to its $k$ nearest neighbors in the infinite crystal structure. Consider an $n\times k$ matrix composed of $n$ rows of distance vectors, where each point $x_{i}\in\mathcal{U}$ corresponds to one row. If the matrix contains $m\geq 1$ identical rows, they are merged into one row with a weight of $\frac{m}{n}$. The resulting matrix can be regarded as a weighted distribution of rows, which is called the Pointwise Distance Distribution $\mathcal{PDD}$$\left(S;k\right)\in\mathbb{R}^{n\times(k+1)}$.
@@
Definition 2: Isometric Crystal Graphs. According to the definition from Widdowson and Kurlin (2022) and Yan et al. (2024a), an isometric transformation is a mapping that preserves Euclidean distances, denoted as $f\left(x\right)=Rx+b$. Any isometric transformation $f$ can be decomposed into translation, rotation, and reflection. Specifically, suppose there exists a rotation matrix $R\in\mathbb{R}^{3\times 3}$, with a determinant of 1 ($\left|R\right|=1$), and a translation vector $b\in\mathbb{R}^{3}$, then two crystal structures $\mathcal{S}=\left(\mathcal{U},\mathcal{L}\right)$ and $\mathcal{Q}=\left(\mathcal{U}^{\prime},\mathcal{L}^{\prime}\right)$ are isometric, satisfying $\mathcal{U}^{\prime}=R\mathcal{U}+b$, where $R\mathcal{U}+b$ denotes the application of the rotation $R$ and translation $b$ to each element in the infinite set $\mathcal{U}$.
@@
If $\mathcal{S}$ and $\mathcal{Q}$ are isometric, then their crystal graph representations satisfy $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$, which means that the graphical representation of the crystal structure produces no false positives; that is, there are no isometric pairs where $\mathcal{G}\left(\mathcal{S}\right)\neq\mathcal{G}\left(\mathcal{Q}\right)$ but $\mathcal{S}\simeq\mathcal{Q}$. Conversely, if $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$, then $\mathcal{S}$ and $\mathcal{Q}$ are isometric, meaning f produces no false negatives, i.e., there are no non-isometric pairs where $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$ but $\mathcal{S}\not\simeq\mathcal{Q}$. That is, if the crystal graph representations of artificially constructed crystal structures are identical under isometric transformations, then they are geometrically equivalent.
@@
Definition 3: Geometrically Complete Crystal Graphs. According to Widdowson and Kurlin (2022) and Yan et al. (2024a), if we construct crystal graphs $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)\Longrightarrow\mathcal{S}\simeq\mathcal{Q}$, where $\simeq$ denotes the isomorphism of two crystals as defined in Definition 2, then the crystal graph $\mathcal{G}$ is geometrically complete. This means that if two crystal graphs $\mathcal{G}\left(\mathcal{S}\right)$ and $\mathcal{G}\left(\mathcal{Q}\right)$ are identical, the infinite crystal structures represented by $\mathcal{G}\left(\mathcal{S}\right)$ and $\mathcal{G}\left(\mathcal{Q}\right)$ are also identical. If the constructed crystal graph $\mathcal{G}$ can distinguish any subtle structural differences between different crystal materials, it is said to be geometrically complete. According to Widdowson and Kurlin (2022), we present Definitions 4-6.
@@
Definition 4: Metric. The metric $d$ between crystal graphs $\mathcal{G}$ satisfies all the axioms: 1) $d\left(\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)\right)=0$ if and only if $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$; 2) Symmetry: $d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)=d\left(\mathcal{G}\left(\mathcal{Q}\right),\mathcal{G}\left(\mathcal{S}\right)\right)$; 3) Triangle inequality: $d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)+d\left(\mathcal{G}\left(\mathcal{Q}\right),\mathcal{G}\left(\mathcal{K}\right)\right)\geq d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{K}\right)\right)$.
@@
Definition 5: Lipschitz continuity of crystal graphs. If $\mathcal{Q}$ is obtained by moving each point in the periodic crystal $\mathcal{S}\subset\mathbb{R}^{n}$ by no more than $\epsilon$, and the distance of the constructed crystal graph structures satisfies $d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{K}\right)\leq C\epsilon\right)$, where $C$ is a constant, then the crystal graph is continuous, and $\mathcal{Q},\mathcal{S}\subset\mathbb{R}^{n}$ can be any periodic crystal structures.
@@
Definition 6: EMD. Let $\mathcal{G}\left(\mathcal{S}\right)$ and $\mathcal{G}\left(\mathcal{Q}\right)$ be the crystal graph structures we construct for periodic crystals $\mathcal{S}$ and $\mathcal{Q}\in\mathbb{R}^{n}$ The flow from $\mathcal{G}\left(\mathcal{S}\right)$ to $\mathcal{G}\left(\mathcal{Q}\right)$ is represented by an $n\left(\mathcal{S}\right)\times n\left(\mathcal{Q}\right)$ matrix, where the elements $f_{ij}\in[0,1]$ indicate the partial flow from $\mathcal{R}_{i}\left(\mathcal{S}\right)$ to $\mathcal{R}_{j}\left(\mathcal{Q}\right)$. The Earth Mover’s Distance (EMD) is defined as the minimum cost: $EMD\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)=\sum_{i=1}^{n}\sum_{j=1}^{n}f_{ij}\left|R_{i}\left(\mathcal{S}\right)-R_{j}\left(\mathcal{Q}\right)\right|$ where $f_{ij}\in[0,1]$ satisfies the following conditions:
$\begin{array}[]{@{}c@{}c@{}}\sum_{i=0}^{n}f_{ij}\leq w_{i}\left(\mathcal{S}\right),\>\sum_{j=0}^{n}f_{ij}\leq w_{j}\left(\mathcal{Q}\right),\>\sum_{i=1}^{n}\sum_{j=1}^{n}f_{ij}=1\end{array}$ (1)
The first condition $\sum_{i=0}^{n}f_{ij}\leq w_{i}\left(\mathcal{S}\right)$ means that not more than the weight $w_{i}\left(\mathcal{S}\right)$ of the component $R_{i}\left(\mathcal{S}\right)$ ‘flows’ into all components $R_{j}\left(\mathcal{Q}\right)$ via ‘flows’$f_{ij}$ Similarly, the second condition $\sum_{j=0}^{n}f_{ij}\leq w_{j}\left(\mathcal{Q}\right)$ means that all ‘flows’ $f_{ij}$ from $R_{i}\left(\mathcal{S}\right)$ ‘flow’ Into $R_{j}\left(\mathcal{Q}\right)$ up to the maximum weight $w_{j}\left(\mathcal{Q}\right)$. The last condition $\sum_{i=1}^{n}\sum_{j=1}^{n}f_{ij}=1$ forces to ‘flow’ all rows $R_{i}\left(\mathcal{S}\right)$ to all rows $R_{j}\left(\mathcal{Q}\right)$.

### v5

@@ Abstract
Crystal structures can be simplified as a periodic point set that repeats across three-dimensional space along an underlying lattice. Traditionally, crystal representation methods characterize the structure using descriptors such as lattice parameters, symmetry, and space groups. However, in reality, atoms in materials always vibrate above absolute zero, causing their positions to fluctuate continuously. This dynamic behavior disrupts the fundamental periodicity of the lattice, making crystal graphs based on static lattice parameters and conventional descriptors discontinuous under slight perturbations. Chemists proposed the pairwise distance distribution (PDD) method to address this problem. However, the completeness of PDD requires defining a large number of neighboring atoms, leading to high computational costs. Additionally, PDD does not account for atomic information, making it challenging to apply it directly to crystal material property prediction tasks. To tackle these challenges, we introduce the atom-Weighted Pairwise Distance Distribution (WPDD) and Unit cell Pairwise Distance Distribution (UPDD) and apply them to the construction of multi-edge crystal graphs. We demonstrate the continuity and general completeness of crystal graphs under slight atomic position perturbations. Moreover, by modeling PDD as global information and integrating it into matrix-based message passing, we significantly reduce computational costs. Comprehensive evaluation results show that WPDDFormer achieves state-of-the-art predictive accuracy across tasks on benchmark datasets such as the Materials Project and JARVIS-DFT.
@@ 1 Introduction
Figure 1: Illustrations of different unit cells and lattice representations of the same crystal structure. Figure (a) shows several possible choices among the infinitely many unit cells for the same crystal structure in the undisturbed case. Figure (b) illustrates that for almost any perturbation, the symmetry group and any reduced unit cell (with minimal volume) will undergo discontinuous changes.
Crystals are solids with a regular geometric shape formed by atoms, ions, or molecules arranged periodically in space during the crystallization process. Their structure is typically described using repeating unit cells and lattice vectors. However, this description method brings a fundamental challenge: the same crystal structure can be represented by different unit cells and lattice vectors, as shown in Figure 1(a). Additionally, in real-world scenarios, the experimental coordinates of unit cells and atoms are inevitably affected by atomic vibrations and measurement noise. These subtle disturbances can lead to discontinuous changes in any simplified unit cell kurlin2024mathematics, resulting in numerous unit cells for a given crystal structure, as shown in Figure 1(b), thereby introducing ambiguity in the representation of crystal data widdowson2022resolving. Currently, many graph neural networks batzner20223; yan2022periodic; yan2024complete; yan2024space typically use unit cell parameters, simplified cell parameters, symmetry, and space groups to represent the periodic structure of crystals. However, these features are either non-invariant or discontinuous zwart2008surprises invariants, leaving the issue of ambiguity in crystal data unresolved patterson1944ambiguities; widdowson2022average; groom2016cambridge; bartok2013representing; wassermann2010activity; ahmad2018machine.
The continuous and complete invariant—Pairwise Distance Distribution (PDD)—proposed by widdowson2022resolving addresses the ambiguity in crystal data representation by distinguishing all periodic structures in the world’s largest real material collection, the Cambridge Structural Database. To achieve completeness, PDD requires a predetermined number of sufficient neighbors, which is computationally expensive and difficult to directly apply for predicting crystal properties balasingham2022compact.  balasingham2024material employed distance distribution graphs (DDGs) based on PDD to predict the properties of crystal materials, but they did not achieve satisfactory performance (only slightly better than CGCNN), and although this approach reduced computational costs, it compromised the completeness of PDD. In contrast, crystal graph representations based on multi-edge crystal graphs and unit cell parameters yan2024complete achieve completeness, more accurately characterizing crystal structures, and achieving state-of-the-art performance in crystal material property prediction tasks. However, using unit cell parameters leads to discontinuities in the crystal graphs.
Since PDD does not account for atomic types, it is challenging to use it directly for effective crystal property prediction. To better represent crystal structures, we first introduce WPDD and UPDD. Furthermore, we integrate WPDD and UPDD into the construction of multi-edge crystal graphs and propose the PDD Graph Transformer (including WPDDFormer and UPDDFormer) based on the transformer architecture. We model WPDD as global information and incorporate it into matrix-based message passing without introducing additional information to capture periodicity, significantly reducing computational costs (as shown in Table 3). Finally, we employ the Earth Mover’s Distance (EMD) rubner2000earth to assess the continuity of crystal graphs, demonstrating that WPDD crystal graphs constructed using only Euclidean distances maintain continuity and general completeness under slight atomic position perturbations, providing a more accurate depiction of actual crystal structures. Ablation experiments show the crucial role of (W/U)PDD in constructing crystal graphs. Through comprehensive evaluations, our method achieves state-of-the-art predictive accuracy across various tasks in the Materials Project and JARVIS datasets. This advancement highlights the effectiveness of WPDDFormer in bridging the gap between traditional crystal descriptors and dynamic atomic behavior, leading to more accurate and reliable predictions in materials science.
@@ 2.1 The Structure of Crystals
By selecting an appropriate structural unit, the entire crystal structure can be viewed as the periodic repetition of this unit in space. This property, where atoms within a crystal repeat in three-dimensional space according to a specific pattern, is called periodicity, with the smallest repeatable structural unit being the unit cell. The unit cell can be defined as $\mathcal{U}=\left(\mathcal{X},\mathcal{P}\right)$, where $\mathcal{X}$ and $\mathcal{P}$ can be represented in matrix form. Typically, $\mathcal{X}=\left[x_{1},x_{2}\cdots x_{n-1},x_{n}\right]^{T}\in\mathbb{R}^{n\times 1}$, where $n$ represents the number of atoms and ${x_{i}}\in\mathbb{R}^{1}$ represents the atomic type of atom i{i} in the unit cell. $\mathcal{P}=\left[p_{1},p_{2}\cdots p_{n-1},p_{n}\right]^{T}\in\mathbb{R}^{n\times 3}$ is the atomic position matrix, where ${p_{i}}\in\mathbb{R}^{3}$ represents the Cartesian coordinates of the atom i{i} in the unit cell in 3D space. The lattice vectors $\mathcal{L}=\left[l_{1},l_{2},l_{3}\right]^{T}\in\mathbb{R}^{3\times 3}$ can reflect the way the unit cell repeats in three directions to map the periodic crystal structure. Therefore, in 3D space, the infinite crystal structure $\mathcal{S}$ can be represented as $\left(\mathcal{U},\mathcal{L}\right)$.
2.2 Definitions
Definition 1.
@@ 2.2 Definitions
Pointwise Distance Distribution. For the infinite crystal structure $\mathcal{S}=\left(\mathcal{U},\mathcal{L}\right)$ mentioned in Section 2.1, fix a neighbor count $k\geq 1$. For each point $x_{i}$ in the unit cell $\mathcal{U}$, let $d_{i1}\leq\dots\leq d_{ik}$ be the Euclidean distances from $\textbf{p}_{i}$ to its $k$ nearest neighbors in the infinite crystal structure. Consider an $n\times k$ matrix composed of $n$ rows of distance vectors, where each point $x_{i}\in\mathcal{U}$ corresponds to one row. If the matrix contains $m\geq 1$ identical rows, they are merged into one row with a weight of $\frac{m}{n}$. The resulting matrix can be regarded as a weighted distribution of rows, which is called the Pairwise Distance Distribution $\mathcal{PDD}$$\left(S;k\right)\in\mathbb{R}^{n\times(k+1)}$.
According to widdowson2022resolving and yan2024complete, we present Definitions 2-3. According to widdowson2022resolving, we present Definitions 4-6.
Definition 2.
@@
Isometric Crystal Graphs. An isometric transformation is a mapping that preserves Euclidean distances, denoted as $f\left(x\right)=Rx+b$. Any isometric transformation $f$ can be decomposed into translation, rotation, and reflection. Specifically, suppose there exists a rotation matrix $R\in\mathbb{R}^{3\times 3}$, with a determinant of 1 ($\left|R\right|=1$), and a translation vector $b\in\mathbb{R}^{3}$, then two crystal structures $\mathcal{S}=\left(\mathcal{U},\mathcal{L}\right)$ and $\mathcal{Q}=\left(\mathcal{U}^{\prime},\mathcal{L}^{\prime}\right)$ are isometric, satisfying $\mathcal{U}^{\prime}=R\mathcal{U}+b$, where $R\mathcal{U}+b$ denotes the application of the rotation $R$ and translation $b$ to each element in the infinite set $\mathcal{U}$.
@@
Suppose $\mathcal{S}$ and $\mathcal{Q}$ are isometric. In that case, their crystal graph representations satisfy $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$, which means that the graphical representation of the crystal structure produces no false positives; that is, there are no isometric pairs where $\mathcal{G}\left(\mathcal{S}\right)\neq\mathcal{G}\left(\mathcal{Q}\right)$ but $\mathcal{S}\simeq\mathcal{Q}$. Conversely, if $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$, then $\mathcal{S}$ and $\mathcal{Q}$ are isometric, meaning f produces no false negatives, i.e., there are no non-isometric pairs where $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$ but $\mathcal{S}\not\simeq\mathcal{Q}$. That is, if the crystal graph representations of artificially constructed crystal structures are identical under isometric transformations, then they are geometrically equivalent.
Definition 3.
@@
Geometrically Complete Crystal Graphs. If we construct crystal graphs $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)\Longrightarrow\mathcal{S}\simeq\mathcal{Q}$, where $\simeq$ denotes the isomorphism of two crystals as defined in Definition 2, then the crystal graph $\mathcal{G}$ is geometrically complete. This means that if two crystal graphs $\mathcal{G}\left(\mathcal{S}\right)$ and $\mathcal{G}\left(\mathcal{Q}\right)$ are identical, the infinite crystal structures represented by $\mathcal{G}\left(\mathcal{S}\right)$ and $\mathcal{G}\left(\mathcal{Q}\right)$ are also identical. If the constructed crystal graph $\mathcal{G}$ can distinguish any subtle structural differences between different crystal materials, it is said to be geometrically complete.
Definition 4.
@@
Metric. The metric $d$ between crystal graphs $\mathcal{G}$ satisfies all the axioms: 1) $d\left(\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)\right)=0$ if and only if $\mathcal{G}\left(\mathcal{S}\right)=\mathcal{G}\left(\mathcal{Q}\right)$; 2) Symmetry: $d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)=d\left(\mathcal{G}\left(\mathcal{Q}\right),\mathcal{G}\left(\mathcal{S}\right)\right)$; 3) Triangle inequality: $d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)+d\left(\mathcal{G}\left(\mathcal{Q}\right),\mathcal{G}\left(\mathcal{K}\right)\right)\geq d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{K}\right)\right)$.
Definition 5.
@@
Lipschitz continuity of crystal graphs. If $\mathcal{Q}$ is obtained by moving each point in the periodic crystal $\mathcal{S}\subset\mathbb{R}^{n}$ by no more than $\epsilon$, and the distance of the constructed crystal graph structures satisfies $d\left(\mathcal{G}\left(\mathcal{S}\right),\mathcal{G}\left(\mathcal{Q}\right)\right)\leq C\epsilon$, where $C$ is a constant, then the crystal graph is continuous, and $\mathcal{Q},\mathcal{S}\subset\mathbb{R}^{n}$ can be any periodic crystal structures.
Definition 6.
@@
EMD. Let $\mathcal{G}\left(\mathcal{S}\right)$ and $\mathcal{G}\left(\mathcal{Q}\right)$ be the crystal graph structures we construct for periodic crystals $\mathcal{S}$ and $\mathcal{Q}\in\mathbb{R}^{n}$ . The flow from $\mathcal{G}\left(\mathcal{S}\right)$ to $\mathcal{G}\left(\mathcal{Q}\right)$ is represented by an $n\left(\mathcal{S}\right)\times n\left(\mathcal{Q}\right)$ matrix, where the elements $f_{ij}\in[0,1]$ indicate the partial flow from $\mathcal{R}_{i}\left(\mathcal{S}\right)$ to $\mathcal{R}_{j}\left(\mathcal{Q}\right)$. The Earth Mover’s Distance (EMD) is defined as the minimum cost:
[…]